What annual rate of return is earned on a $5,000 investment when it grows to $10,250 in five years? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Please show step by step
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

10250=5000*(1+r/100)^5

(10250/5000)^(1/5)=(1+r/100)

(1+r/100)=1.1544

r=1.1544-1

=15.44%(Approx).